### v1

# On the Interplay between Human Label Variation and Model Fairness

Abstract
The impact of human label variation (HLV) on model fairness is an unexplored topic. This paper examines the interplay by comparing training on majority-vote labels with a range of HLV methods. Our experiments show that without explicit debiasing, HLV training methods have a positive impact on fairness.111Code and supplementary materials: https://github.com/kmkurn/hlv-fairness-interplay
On the Interplay between Human Label Variation and Model Fairness
Kemal Kurniawan1 Meladel Mistica1 Timothy Baldwin2,1 Jey Han Lau1 1School of Computing and Information Systems, University of Melbourne 2Mohamed bin Zayed University of Artificial Intelligence {kurniawan.k,misticam,laujh}@unimelb.edu.au tb@ldwin.net
1 Introduction
Human label variation (HLV; Plank, 2022) has been shown to improve performance through better generalisation (Peterson et al., 2019; Uma et al., 2021b; Kurniawan et al., 2025). However, the impact of HLV on model fairness remains unexplored. It is well established that there is a trade-off between performance and fairness (Han et al., 2022; Shen et al., 2022), so one might conjecture that HLV would similarly negatively impact fairness (since HLV improves performance). We explore this interplay between HLV and model fairness in this paper, in what we believe to be the first systematic analysis of this interaction. We compare training on majority-vote labels with four HLV training methods from prior work (Kurniawan et al., 2025), and evaluate on two tasks: offensiveness classification with SBIC (Sap et al., 2020) and legal area classification with TAG, an in-house legal dataset.
We find that across SBIC and TAG, training with HLV methods improves overall performance without harming fairness. Over TAG, training with HLV even improves fairness. Since these findings are only from a single set of class and group weights (configuration hereafter), we examine many possible configurations to draw more robust conclusions. For both datasets, HLV does not reduce model fairness for most configurations. Interestingly, HLV can improve model fairness for a substantial number of configurations.
We then investigate why HLV improves fairness by introducing temperature scaling to an HLV training method to control the weighting of minority annotations. Our TAG experiments suggest fairness gains indeed come from minority annotations.
Our findings suggest that training with HLV improves performance without compromising fairness. That said, our findings also underscore the effect of class and group weightings in measuring fairness in the HLV context. Therefore, it is important to understand what fairness means for each application before incorporating HLV.
2 Related Work
HLV (Plank, 2022) challenges the standard assumption that an instance has one ground truth, in line with data perspectivism (Cabitza et al., 2023) which advocates for the consideration of multiple perspectives in data annotation. Previous work on this topic includes model generalisation (Uma et al., 2021b, a; Leonardelli et al., 2023), evaluation metrics (Uma et al., 2021b; Rizzi et al., 2024; Kurniawan et al., 2025), model calibration (Baan et al., 2022), and separating annotation signal from noise (Weber-Genzel et al., 2024; Ivey et al., 2025). HLV has also been shown to improve model generalisation for computer vision tasks (Peterson et al., 2019). Our work contributes to this growing discipline by examining the interplay between HLV and fairness.
Previous work covers the interplay between model fairness and other aspects of machine learning models. Hessenthaler et al. (2022) studied how model fairness and environmental costs affect each other. Zhao et al. (2023); Brandl et al. (2024) explored the interaction between model fairness and explainability, as did Ferry et al. (2025) but with privacy considerations. In contrast, we study how model fairness is impacted by HLV, which to our knowledge, is the first examination of this interplay.
3 Method
3.1 Datasets
SBIC
Each instance in the Social Bias Inference Corpus (Sap et al., 2020) is an English social media post annotated with an offensiveness class and a list of target minority groups.222Offensiveness classes = Yes/No/Maybe; example target groups = black folks, asian folks, etc. Target groups are divided into 7 broad categories333Body, culture, disabled, gender, race, social, victim. with each post annotated by 3 crowd annotators. We compute its offensiveness class distribution and union of the broad categories over these 3 annotations.
TAG
[…]
3.2 Evaluation
To evaluate fairness, we first compute a fairness score (defined later) for each class and group . Next, following Han et al. (2023), we aggregate over groups then classes using weighted generalised mean to obtain the overall fairness score. The group-wise aggregation is defined as
where is a scalar exponent controlling the contribution of smaller-valued and is a non-negative importance weights for group satisfying . The class-wise aggregation is performed analogously over using separate class-specific weights and exponent.
To define the groups, we use the broad target categories for SBIC, and the defined cohorts for TAG. Therefore, for SBIC, the group information is associated with the target of the social media post, while the group information for TAG is associated with the help-seeker who produced the instance.
In both datasets, each instance can belong to zero or more groups. Therefore, we partition the instances into two subsets for each group based on in- or out-of-group membership, respectively.777See Appendix B for subset size statistics. We compute by first evaluating performance for class on both subsets then taking the ratio between the worse- and the better-performing subsets, following Yeh et al. (2024).888See Section C.1 for a more formal definition.
Computing requires a performance metric for each class that works in the HLV context (i.e., soft ground truths). Thus, we use soft F1, the only metric satisfying the requirement (Kurniawan et al., 2025). For a given class , it measures the amount of overlap between the ground-truth and predicted probabilities of class relative to their total.999See Section C.2 for details.
Aggregation configurations
For TAG results in Section˜4.1, the class and group aggregation weights are provided by the legal organisation who co-developed the dataset (Section˜3.1). For SBIC, as we have no a priori preference we weight each class and group equally. For both datasets, we set for both aggregation, resulting in a standard weighted average. In Section˜4.2, we experiment with multiple randomly-sampled configurations for more robust conclusions.101010Details are given in Appendix E.
Overall performance
To assess performance in the HLV context, we use soft micro F1, a natural extension of class-wise soft F1 which correlates well with human judgement (Kurniawan et al., 2025).111111See Section C.2 for details.
3.3 Approaches
We experiment with two pretrained models: base RoBERTa (Liu et al., 2019, 100M parameters) and Nov’24 7B OLMo 2 (Walsh et al., 2025). We replace the output layer of each model with the appropriate classification layer121212A linear layer then either a softmax (SBIC) or sigmoid (TAG) activation functions. and finetune on the train set.
For HLV training methods, we use the four most successful methods identified by Kurniawan et al. (2025): repeated labelling (ReL), soft labelling (SL), minimising Jensen-Shannon divergence (JSD), and maximising soft micro F1 score (SmF1).131313See Appendix F for more details. For non-HLV baselines, we finetune them on majority-voted classes (MV).141414Further training details are reported in Appendix G.
4 Results
4.1 Single Fairness Configuration
We report the overall performance and fairness of each method in Table˜1. For SBIC, the table shows that HLV training methods consistently outperform MV for overall performance across the two models. These improvements are statistically significant (with bootstrap tests) for 6 out of 8 model–method pairs.151515Exceptions: SL and JSD with RoBERTa As for fairness, the HLV training methods can improve or reduce fairness, but the differences are not statistically significant for 7 out of 8 model–method pairs.161616Exception: SmF1 with RoBERTa
Looking at the TAG results, for both overall performance and fairness, the HLV training methods consistently outperform MV across both models, with statistically significant improvements for 6 out of 8 model–method pairs.171717Exception: SmF1 Taken together, the results suggest that HLV training improves overall performance without sacrificing fairness, and in some cases (e.g., TAG), can even improve fairness.
| Model | Method | SBIC | TAG | ||
| Perf | Fair | Perf | Fair | ||
| OLMo | MV | ||||
| ReL | ∗ | ∗ | ∗ | ||
| SL | ∗ | ∗ | ∗ | ||
| JSD | ∗ | ∗ | ∗ | ||
| SmF1 | ∗ | ||||
| RoBERTa | MV | ||||
| ReL | ∗ | ∗ | ∗ | ||
| SL | ∗ | ∗ | |||
| JSD | ∗ | ∗ | |||
| SmF1 | ∗ | ∗ | |||
4.2 Multiple Fairness Configurations
The results in Section˜4.1 were obtained with one configuration, i.e. fixed values for and in Equation˜1 for both group- and class-wise aggregation. Here, we test multiple randomly-sampled configurations to draw more robust conclusions.181818See Appendix E for more details. This experimental setup is especially important for SBIC: the single configuration used in Section˜4.1 is divorced from any intended application of the model.
For each randomly-sampled configuration,191919Note that performance is constant regardless of configurations as they configure only the fairness score computation. we perform a two-tailed bootstrap test across 3 runs to check if the mean fairness score difference between each HLV training method and MV is statistically significant (). If so, we record if the mean fairness score of MV is higher (i.e. fairer). Otherwise, we deem the HLV training method as fair as MV. We report the fraction of configurations where MV is found not fairer than each HLV method on SBIC in Figure˜1. TAG follows this trend.202020The comparisons are shown in Appendix H.
[…]
That said, Figure˜1 also shows a substantial number of configurations where MV is fairer than training with HLV. This finding highlights the need to determine fairness configurations specific to the application before incorporating HLV.
5 Analysis
5.1 Fairness Configurations
[…]
Scalar controls the contribution of smaller-valued in group-wise (Equation˜1) and class-wise aggregation. Smaller/larger means prioritising lower-scoring groups or classes more/less. Therefore, the results suggest that when one cares about all classes or groups similarly regardless of their scores (mid), training with HLV does not reduce fairness. This setting corresponds to most existing fairness evaluation (Han et al., 2023). However, when one prioritises the lowest-scoring classes or groups in fairness (low, similar to Rawls (2001)), training with HLV is comparable to MV.212121The results also suggest that training with HLV generally never harms fairness when one focuses more on the highest-scoring classes or groups (high). However, this overestimates model fairness and seems unused in practice.
5.2 Minority Annotations
Here, we investigate factors that make HLV beneficial for model fairness. We focus on the TAG dataset on which HLV generally outperforms MV in both performance and fairness as reported in Section˜4.1. We hypothesise that the fairness improvements come from the minority annotations.
To test this, we introduce temperature scaling to SL. Each class probability is proportional to the number of annotators that select the class raised to the power of where is the temperature. As , the most-voted class will have probability close to 1, resembling MV. In contrast, the class distribution will reduce to uniform as , thereby weighting up the minority annotations.222222In practice, distributions will be uniform over all most-voted classes as and over selected classes as . Standard SL corresponds to .
Figure˜3 shows when is small, performance and fairness are low. As increases, performance gradually climbs but peaks at , then declines as grows further. In contrast, fairness improves as grows, then stabilises after peaking at . These findings suggest minority annotations contain useful signals for fairness.232323We offer a possible explanation in Appendix J. Unsurprisingly, results show performance peaks at because the scaling is only applied to the train set, i.e. test ground truths are computed normally ().
6 Conclusions
While HLV has been shown to improve generalisation, this is the first work to study the interplay between HLV and fairness. We first showed empirically that training with HLV does not harm model fairness, and indeed, it can improve fairness. We then showed that this observation is robust to various fairness configurations. However, when the lowest-scoring groups or classes in fairness were prioritised, training with HLV provided no fairness benefits compared to training on majority-voted classes. Lastly, we confirmed that the fairness improvements with HLV training come from the minority annotations. Our findings highlight the performance and fairness benefits of training with HLV, but we recommend NLP practitioners to consider what fairness means in their application before incorporating HLV.
Limitations
Because the TAG dataset contains real-world confidential legal help requests, we cannot distribute it as it used in the study. However, we still include it because of its importance to our analysis in Section˜5.2, and we believe that the insight obtained from the dataset in the study is still useful. Although we cannot distribute the dataset, we can distribute ground-truth and predicted judgement distributions of the dev and the test sets.242424https://github.com/kmkurn/hlv-fairness-interplay This will allow other researchers to replicate our reported numbers and build on our work to some extent.
We evaluate using soft evaluation metrics in all experiments. One could argue that this setup favours HLV training, so the comparisons with MV are unfair. However, we note that previous HLV work has found that MV outperforms many HLV training methods in this setup (Kurniawan et al., 2025), suggesting that the argument doesn’t hold in practice.
Another limitation of our study is the exclusive use of English datasets. This is because of the scarcity of datasets that have both disaggregated annotations for HLV and instance attributes (e.g., cohort information) which are appropriate for fairness evaluation. However, we expect that our findings would generalise to non-English languages, as our experimental design is language-independent. We encourage future work to develop such datasets for non-English languages to test this empirically.
Acknowledgments
This research is supported by the Australian Research Council Linkage Project LP210200917252525https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government. This research is done in collaboration with Justice Connect, an Australian public benevolent institution.262626As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project. This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
Appendix A TAG Dataset Details
The TAG dataset has a total of 11K instances split randomly 8:1:1 for train, dev, and test sets. Each instance is a legal problem description, e.g., my landlord evicted me w/o notice, annotated by an average of 5.5 lawyers with one or more legal areas out of 33 options, e.g., Criminal law, Elder law, Not a legal issue. Each instance is also associated with one or more cohorts out of 6 choices, e.g., seniors, low income earners. Sensitive identifying information had been anonymised by the owner organisation prior to the dataset use in our work.
Appendix B Subset Size Statistics
Table˜2 reports the number of instances in the in- and out-of-group subsets for each group in the datasets. For TAG, the help-seekers were allowed to select “Prefer not to say” when self-reporting. For fairness evaluation, we include only those that did not select that option, i.e. their identities were given explicitly.
| Group | Subset | |
| In-group | Out-of-group | |
| Body | 58 | 4640 |
| Culture | 495 | 4203 |
| Disabled | 112 | 4586 |
| Gender | 503 | 4195 |
| Race | 819 | 3879 |
| Social | 104 | 4594 |
| Victim | 215 | 4483 |
| Group | Subset | |
| In-group | Out-of-group | |
| ATS | 35 | 317 |
| HOM | 154 | 64 |
| LGB | 29 | 8 |
| LOW | 135 | 148 |
| PUB | 19 | 109 |
| SEN | 164 | 90 |
Appendix C Evaluation Details
C.1 Definition of
If and denote the performance for class on the in- and out-of-group subsets defined by group respectively, then the fairness score is defined as
C.2 Performance
To compute and in Equation˜2, we use the soft F1 score (Kurniawan et al., 2025). The score for class is defined as
where and denote the ground-truth and the predicted probability of class for instance respectively.
To evaluate overall performance with soft ground truths, we use the soft micro F1 score (Kurniawan et al., 2025). It is defined formally as
This equation is similar to Equation˜3, but the summations are performed over both instances and classes.
Appendix D Difference in Fairness Interpretations
Due to the different group definitions for both datasets, the interpretations of fairness are also different. Consider an arbitrary group in SBIC. Here, a fair model is one that performs similarly in identifying offensiveness on social media posts that target and do not target the minority category corresponding to the group. This is sensible because we do not want a model to identify offensive posts less accurately when the posts target minorities. In contrast, for an arbitrary group in TAG, a fair model is one that achieves similar levels of performance for help-seekers that belong and do not belong to the group. This is also sensible because we do not want a model to more frequently fail at recognising relevant legal areas for a particular cohort (e.g., seniors). Thus, while there is a difference, the fairness definitions are still reasonable for both cases.
Appendix E Configuration Sampling
To get the results in Section˜4.2, we randomly sample ten thousand configurations for each of group- and class-wise aggregation. Each configuration consists of a set of item weights and a scalar exponent that parameterises the weighted generalised mean used for aggregation. Specifically, the weights and exponent are drawn from a flat Dirichlet distribution and respectively. Smaller and larger values of make the generalised mean focus more on lower and higher fairness scores respectively. In particular, as and , generalised mean reduces to and respectively. In other words, the exponent controls how much the lowest-scoring groups or classes are prioritised.
Appendix F HLV Training Methods
We experiment with four HLV training methods:
1.
ReL (repeated labelling) which treats each annotation of an instance as a separate instance–class pair. In other words, an instance may appear multiple times in the training data, each with a different class. The loss function is standard cross-entropy.
2.
SL (soft labelling) which computes a class distribution for each instance and uses the distributions as soft ground truths for training. Each class probability is proportional to the number of annotators that select the class. Standard cross-entropy is used as the loss function.
3.
JSD which computes class distributions similarly to SL but uses Jensen-Shannon divergence as the loss function.
4.
SmF1 which also computes class distributions as SL does but maximises the soft micro F1 score during training (see Section˜C.2).
We refer the readers to the work by Kurniawan et al. (2025) for more details on these methods.
Appendix G Training Details
We implement all methods using FlairNLP (Akbik et al., 2019).272727https://flairnlp.github.io For RoBERTa, we tune the learning rate and the batch size using random search, optimising only for overall performance (without fairness). Table˜3 reports the best values found. For OLMo, we use FlairNLP’s default hyperparameters282828Learning rate and batch size are and 32 respectively. for computational reasons. We use LoRA (Hu et al., 2022) to finetune OLMo efficiently. All models are finetuned for 10 epochs. On a single NVIDIA A100 GPU with 80GB of memory, the finetuning runs for:
less than an hour with RoBERTa for all methods except ReL on SBIC and TAG,
slightly over 2 hours with RoBERTa for ReL on SBIC and TAG,
slightly over 10 hours with OLMo for all methods except ReL on SBIC,
about 1.3 days with OLMo for ReL on SBIC,
about 9 hours with OLMo for all methods except ReL on TAG, and
about 2 days with OLMo for ReL on TAG.
We truncate inputs longer than 512 tokens in TAG to reduce memory consumption, affecting only 4% of instances.
| Dataset | Method | Learning rate | Batch size |
| SBIC | MV | 4 | |
| ReL | 4 | ||
| SL | 4 | ||
| JSD | 512 | ||
| SmF1 | 512 | ||
| TAG | MV | 16 | |
| ReL | 8 | ||
| SL | 8 | ||
| JSD | 16 | ||
| SmF1 | 64 |
Appendix H More Multiple Fairness Configurations Results
Figure˜4 shows the fraction of configurations where MV is not fairer than each HLV training method on TAG. The figure shows a similar trend to that of SBIC reported in Section˜4.2.
[…]
Appendix I More Fairness Configurations Analysis Results
This section shows the fraction of configurations where MV is not fairer than each HLV training method for various levels of exponent . Figure˜6 shows the results on SBIC for the exponent in the class-wise aggregation (the group-wise counterpart is shown in Figure˜2). In contrast, Figure˜7 shows the same results on TAG for the exponent in both the group- and class-wise aggregation. The figures show that the trend is similar to that reported in Section˜5.1.
Appendix J Possible Explanation on Minority Annotations Improving Fairness on TAG
We offer a possible explanation for the finding in Section˜5.2 that the minority annotations on TAG contain useful signals that improve fairness. TAG was annotated by lawyers, each with a specific set of legal areas of specialisation. Intuitively, some legal specialisations are more common than others. We hypothesise that minority annotations are correlated with rare specialisations, which also correlate with certain cohorts. This would imply that incorporating the minority annotations would result in fairer models for the defined cohorts.

### v2

@@
 This paper examines the interplay by comparing training on majority-vote labels with a range of HLV methods.
-Our experiments show that without explicit debiasing, HLV training methods have a positive impact on fairness.111Code and supplementary materials: https://github.com/kmkurn/hlv-fairness-interplay
+Our experiments show that without explicit debiasing, HLV training methods have a positive impact on fairness under certain configurations.111Code and supplementary materials: https://github.com/kmkurn/hlv-fairness-interplay
 On the Interplay between Human Label Variation and Model Fairness
@@
 However, the impact of HLV on model fairness remains unexplored.
-It is well established that there is a trade-off between performance and fairness (Han et al., 2022; Shen et al., 2022), so one might conjecture that HLV would similarly negatively impact fairness (since HLV improves performance).
+One might conjecture that HLV would improve fairness because it preserves minority views and avoids collapsing them into a single point like majority voting.
+However, it is well established that there is a trade-off between performance and fairness (Han et al., 2022; Shen et al., 2022), so one might also conjecture that HLV would similarly negatively impact fairness (since HLV improves performance).
 We explore this interplay between HLV and model fairness in this paper, in what we believe to be the first systematic analysis of this interaction.
-We compare training on majority-vote labels with four HLV training methods from prior work (Kurniawan et al., 2025), and evaluate on two tasks: offensiveness classification with SBIC (Sap et al., 2020) and legal area classification with TAG, an in-house legal dataset.
+We compare training on majority-vote labels with four HLV training methods from prior work (Kurniawan et al., 2025), and evaluate on two tasks: offensiveness classification with SBIC (Sap et al., 2020) and legal area classification with TAG, an in-house legal dataset, developed with practising lawyers.
 We find that across SBIC and TAG, training with HLV methods improves overall performance without harming fairness.
@@
 Since these findings are only from a single set of class and group weights (configuration hereafter), we examine many possible configurations to draw more robust conclusions.
-For both datasets, HLV does not reduce model fairness for most configurations.
+For both datasets, HLV does not reduce model fairness for most configurations, and this trend is clearer when the configuration prioritises all groups/classes equally.
 Interestingly, HLV can improve model fairness for a substantial number of configurations.
 We then investigate why HLV improves fairness by introducing temperature scaling to an HLV training method to control the weighting of minority annotations.
-Our TAG experiments suggest fairness gains indeed come from minority annotations.
+Our TAG experiments suggest fairness gains do indeed come from minority annotations.
 Our findings suggest that training with HLV improves performance without compromising fairness.
-That said, our findings also underscore the effect of class and group weightings in measuring fairness in the HLV context.
+Additionally, our findings underscore the effect of class and group weightings in measuring fairness in the HLV context.
 Therefore, it is important to understand what fairness means for each application before incorporating HLV.
@@
 3.2 Evaluation
-To evaluate fairness, we first compute a fairness score (defined later) for each class and group .
+We focus on a group-level parity definition of fairness in this study due to its prevalence in prior work (Shen et al., 2022; Han et al., 2023).
+To evaluate this, we first compute a fairness score (defined later) for each class and group .
 Next, following Han et al.
@@
 For SBIC, as we have no a priori preference we weight each class and group equally.
-For both datasets, we set for both aggregation, resulting in a standard weighted average.
+For both datasets, we set for both aggregations, resulting in standard weighted average.
 In Section˜4.2, we experiment with multiple randomly-sampled configurations for more robust conclusions.101010Details are given in Appendix E.
@@
 (2025): repeated labelling (ReL), soft labelling (SL), minimising Jensen-Shannon divergence (JSD), and maximising soft micro F1 score (SmF1).131313See Appendix F for more details.
-For non-HLV baselines, we finetune them on majority-voted classes (MV).141414Further training details are reported in Appendix G.
+For non-HLV baselines, we finetune them on majority-vote classes (MV).141414Further training details are reported in Appendix G.
 4 Results
@@
 For SBIC, the table shows that HLV training methods consistently outperform MV for overall performance across the two models.
-These improvements are statistically significant (with bootstrap tests) for 6 out of 8 model–method pairs.151515Exceptions: SL and JSD with RoBERTa As for fairness, the HLV training methods can improve or reduce fairness, but the differences are not statistically significant for 7 out of 8 model–method pairs.161616Exception: SmF1 with RoBERTa
-Looking at the TAG results, for both overall performance and fairness, the HLV training methods consistently outperform MV across both models, with statistically significant improvements for 6 out of 8 model–method pairs.171717Exception: SmF1 Taken together, the results suggest that HLV training improves overall performance without sacrificing fairness, and in some cases (e.g., TAG), can even improve fairness.
+These improvements are statistically151515We report effect sizes as a measure of practical significance in Appendix H.
+significant (with bootstrap tests) for 6 out of 8 model–method pairs.161616Exceptions: SL and JSD with RoBERTa As for fairness, the HLV training methods can improve or reduce fairness, but the differences are not statistically significant for 7 out of 8 model–method pairs.171717Exception: SmF1 with RoBERTa
+Looking at the TAG results, for both overall performance and fairness, the HLV training methods consistently outperform MV across both models, with statistically significant improvements for 6 out of 8 model–method pairs.181818Exception: SmF1 Taken together, the results suggest that HLV training improves overall performance without sacrificing fairness, and in some cases (e.g., TAG), can even improve fairness.
+Discussion
+Unsuprisingly, the different HLV methods attain different levels of fairness.
+However, these variations are consistent with their overall performance.
+For example, ReL, a strong method performance-wise, is also among the strongest in terms of fairness, while SmF1 is among the weakest.
+That SmF1 attains poor performance is not surprising because Kurniawan et al.
+(2025) showed that the method can result in degenerate predictions where all probability mass is concentrated on the majority class.
+Looking into class- and group-wise performance for each method, we find that all noticeable differences seem to align well with overall performance trends.
 | Model | Method | SBIC | TAG | ||
@@
 fixed values for and in Equation˜1 for both group- and class-wise aggregation.
-Here, we test multiple randomly-sampled configurations to draw more robust conclusions.181818See Appendix E for more details.
+Here, we test multiple randomly-sampled configurations to draw more robust conclusions.191919See Appendix E for more details.
 This experimental setup is especially important for SBIC: the single configuration used in Section˜4.1 is divorced from any intended application of the model.
-For each randomly-sampled configuration,191919Note that performance is constant regardless of configurations as they configure only the fairness score computation.
+For each randomly-sampled configuration,202020Note that performance is constant regardless of configurations as they configure only the fairness score computation.
 we perform a two-tailed bootstrap test across 3 runs to check if the mean fairness score difference between each HLV training method and MV is statistically significant ().
@@
 Otherwise, we deem the HLV training method as fair as MV.
-We report the fraction of configurations where MV is found not fairer than each HLV method on SBIC in Figure˜1.
-TAG follows this trend.202020The comparisons are shown in Appendix H.
+Note that overall performance is unaffected by fairness configurations.
+Therefore, we compare only the fairness scores.
+We report the fraction of configurations where MV is found not to be fairer than each HLV method on SBIC in Figure˜1. The result on TAG follows the same trend, which we show in Appendix˜I.
 […]
@@
 Standard SL corresponds to .
-Figure˜3 shows when is small, performance and fairness are low.
-As increases, performance gradually climbs but peaks at , then declines as grows further.
-In contrast, fairness improves as grows, then stabilises after peaking at .
-These findings suggest minority annotations contain useful signals for fairness.232323We offer a possible explanation in Appendix J.
-Unsurprisingly, results show performance peaks at because the scaling is only applied to the train set, i.e.
-test ground truths are computed normally ().
+We report how fairness changes as grows in Figure˜3.
+The figure shows that when is small, fairness is low.
+As increases, fairness improves and stabilises after peaking at .
+These findings suggest minority annotations contain useful signals for fairness.
+We offer a possible explanation for this in Appendix˜K and leave the full investigation on this direction for future work.
+Unsurprisingly, the figure also shows that performance peaks at . This is because the scaling is only applied to the train set, i.e. test ground truths are computed normally ().
 6 Conclusions
@@
 We then showed that this observation is robust to various fairness configurations.
-However, when the lowest-scoring groups or classes in fairness were prioritised, training with HLV provided no fairness benefits compared to training on majority-voted classes.
+However, when the lowest-scoring groups or classes in fairness were prioritised, training with HLV provided no fairness benefits compared to training on majority-vote classes.
 Lastly, we confirmed that the fairness improvements with HLV training come from the minority annotations.
-Our findings highlight the performance and fairness benefits of training with HLV, but we recommend NLP practitioners to consider what fairness means in their application before incorporating HLV.
+Our findings highlight the performance and fairness benefits of training with HLV under certain configurations.
+Therefore, we recommend NLP practitioners to: (1) consider what fairness means in their application, (2) define it in terms of the fairness evaluation configuration described in Section˜3.2, (3) incorporate HLV if the configuration prioritises all groups/classes roughly equally, and (4) evaluate fairness with the configuration to confirm that HLV does not harm fairness compared to majority voting.
 Limitations
@@
 However, we still include it because of its importance to our analysis in Section˜5.2, and we believe that the insight obtained from the dataset in the study is still useful.
-Although we cannot distribute the dataset, we can distribute ground-truth and predicted judgement distributions of the dev and the test sets.242424https://github.com/kmkurn/hlv-fairness-interplay This will allow other researchers to replicate our reported numbers and build on our work to some extent.
+Although we cannot distribute the dataset, we can distribute ground-truth and predicted judgement distributions of the dev and the test sets.232323https://github.com/kmkurn/hlv-fairness-interplay This will allow other researchers to replicate our reported numbers and build on our work to some extent.
+Prior work has evaluated fairness using synthetic, targeted check lists (Manerba and Tonelli, 2021).
+Therefore, a reasonable solution may be to follow a similar approach to address the data privacy issue.
+We agree in theory that it would allow us to not only measure fairness at a high level but also understand in what cases the models break down.
+However, our goal is to understand fairness and its relationship with HLV rather than address specific impacts of model bias.
+Furthermore, we foresee a few complications: (1) it is not always immediately clear what the biases might be as we need to know them in order to create the synthetic check lists); (2) in our case, the language of the TAG data is non-standard English, and so it is difficult to mimic that genre; and (3) we will need to distribute our TAG-trained models, which is restricted because of IP constraints and possible risks of leaking the sensitive TAG training data through the model.
+Therefore, this work does not take this approach.
 We evaluate using soft evaluation metrics in all experiments.
 One could argue that this setup favours HLV training, so the comparisons with MV are unfair.
-However, we note that previous HLV work has found that MV outperforms many HLV training methods in this setup (Kurniawan et al., 2025), suggesting that the argument doesn’t hold in practice.
+However, we note that previous work has found that MV outperforms many HLV training methods in this setup (Kurniawan et al., 2025), suggesting that the argument does not hold in practice.
 Another limitation of our study is the exclusive use of English datasets.
@@
 Acknowledgments
-This research is supported by the Australian Research Council Linkage Project LP210200917252525https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government.
-This research is done in collaboration with Justice Connect, an Australian public benevolent institution.262626As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project.
+This research is supported by the Australian Research Council Linkage Project LP210200917242424https://dataportal.arc.gov.au/NCGP/Web/Grant/Grant/LP210200917 and funded by the Australian Government.
+This research is done in collaboration with Justice Connect, an Australian public benevolent institution.252525As defined by the Australian government: https://www.acnc.gov.au/charity/charities/4a24f21a-38af-e811-a95e-000d3ad24c60/profile We thank Kate Fazio, Tom O’Doherty, and Rose Hyland from Justice Connect for their support throughout the project.
 This research is supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
@@
 Appendix G Training Details
-We implement all methods using FlairNLP (Akbik et al., 2019).272727https://flairnlp.github.io For RoBERTa, we tune the learning rate and the batch size using random search, optimising only for overall performance (without fairness).
+We implement all methods using FlairNLP (Akbik et al., 2019).262626https://flairnlp.github.io For RoBERTa, we tune the learning rate and the batch size using random search, optimising only for overall performance (without fairness).
 Table˜3 reports the best values found.
-For OLMo, we use FlairNLP’s default hyperparameters282828Learning rate and batch size are and 32 respectively.
+For OLMo, we use FlairNLP’s default hyperparameters272727Learning rate and batch size are and 32 respectively.
 for computational reasons.
@@
 | SmF1 | 64 |
-Appendix H More Multiple Fairness Configurations Results
+Appendix H Effect Sizes
+[…]
+| Model | Method | SBIC | TAG | ||||||
+| Perf | Fair | Perf | Fair | ||||||
+| low | high | low | high | low | high | low | high | ||
+| OLMo | ReL | ||||||||
+| SL | |||||||||
+| JSD | |||||||||
+| SmF1 | |||||||||
+| RoBERTa | ReL | ||||||||
+| SL | |||||||||
+| JSD | |||||||||
+| SmF1 | |||||||||
+Appendix I More Multiple Fairness Configurations Results
 Figure˜4 shows the fraction of configurations where MV is not fairer than each HLV training method on TAG. The figure shows a similar trend to that of SBIC reported in Section˜4.2.
 […]
-Appendix I More Fairness Configurations Analysis Results
+Appendix J More Fairness Configurations Analysis Results
 This section shows the fraction of configurations where MV is not fairer than each HLV training method for various levels of exponent .
@@
 The figures show that the trend is similar to that reported in Section˜5.1.
-Appendix J Possible Explanation on Minority Annotations Improving Fairness on TAG
+Appendix K Possible Explanation on Minority Annotations Improving Fairness on TAG
 We offer a possible explanation for the finding in Section˜5.2 that the minority annotations on TAG contain useful signals that improve fairness.